Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A magician has five animals in his magic hat: 3 doves and 2

Expert replies
by AAPL » Sat Feb 16, 2019 3:16 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

GMAT Prep

A magician has five animals in his magic hat: 3 doves and 2 rabbits. If he pulls two animals out of the hat at random, what is the chance that he will have a matched pair?

A. 2/5
B. 3/5
C. 1/5
D. 1/2
E. 7/5

OA A
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Sat Feb 16, 2019 3:45 am
AAPL wrote:GMAT Prep

A magician has five animals in his magic hat: 3 doves and 2 rabbits. If he pulls two animals out of the hat at random, what is the chance that he will have a matched pair?

A. 2/5
B. 3/5
C. 1/5
D. 1/2
E. 7/5
AND means MULTIPLY.
OR means ADD.

Case 1: 2 doves
P(1st animal is a dove) = 3/5. (Of the 5 animals, 3 are doves.)
P(2nd animal is a dove) = 2/4. (Of the 4 remaining animals, 2 are doves.)
Since in Case 1 we need the 1st animal to be a dove AND the 2nd animal to be a dove, we MULTIPLY the fractions:
3/5 * 2/4 = 6/20 = 3/10

Case 2: 2 rabbits
P(1st animal is a rabbit) = 2/5. (Of the 5 animals, 2 are rabbits.)
P(2nd animal is a rabbit) = 1/4. (Of the 4 remaining animals, 1 is a rabbit.)
Since in Case 2 we need the 1st animal to be a rabbit AND the 2nd animal to be a rabbit, we MULTIPLY the fractions:
2/5 * 1/4 = 2/20 = 1/10

Since a good outcome will be yielded by Case 1 OR by Case 2, we ADD the blue fractions:
3/10 + 1/10 = 4/10 = 2/5

The correct answer is A.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Brent@GMATPrepNow » Sat Feb 16, 2019 6:30 am
AAPL wrote:GMAT Prep

A magician has five animals in his magic hat: 3 doves and 2 rabbits. If he pulls two animals out of the hat at random, what is the chance that he will have a matched pair?

A. 2/5
B. 3/5
C. 1/5
D. 1/2
E. 7/5

OA A
Let's apply probability rules

First notice that, to get a matched pair, we can select 2 doves or 2 rabbits.

So, P(matched pair) = P(1st pick is rabbit AND 2nd pick is rabbit OR 1st pick is dove AND 2nd pick is dove)

We can now apply our AND and OR rules to get:
P(matched pair) = [P(1st pick is rabbit) x P(2nd pick is rabbit)] + [P(1st pick is dove) x P(2nd pick is dove)]
So, P(matched pair) = (3/5 x 2/4) + (2/5 x 1/4)
We get: 2/5 (or 0.4)

Answer: A


Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Fri Feb 22, 2019 3:18 pm
AAPL wrote:GMAT Prep

A magician has five animals in his magic hat: 3 doves and 2 rabbits. If he pulls two animals out of the hat at random, what is the chance that he will have a matched pair?

A. 2/5
B. 3/5
C. 1/5
D. 1/2
E. 7/5

OA A
We are given that from a group of 3 doves and 2 rabbits, 2 animals will be randomly selected. We need to determine the probability that a matched pair will be pulled out of the hat.

In other words, we need to determine:

P(2 doves pulled) + P(2 rabbits pulled)

We can use combinations to determine the number of favorable outcomes (that 2 rabbits or 2 doves are selected) and the total number of outcomes (that 2 animals are selected from 5).

Let’s first determine the number of ways we can select 2 doves from 3:

# of ways to select 2 doves from 3 doves: 3C2 = 3

Next let’s determine the number of ways we can select 2 rabbits from 2:

# of ways to select 2 rabbits from 2 rabbits: 2C2 = 1

Now we must determine the number of ways to select 2 animals from a total of 5 animals:

5C2 = (5 x 4)/(2 x 1) = 10

Thus, the probability of selecting a matched pair is 3/10 + 1/10 = 4/10 = 2/5.

Alternate Solution:

The two events that satisfy the requirement of getting a matched pair are DD or RR.

The probability of DD is 3/5 x 2/4 = 6/20 =3/10.

The probability of RR is 2/5 x 1/4 = 2/20 = 1/10.

Since DD and RR are mutually exclusive events, the probability that either of these two events happens can be found simply by adding the individual probabilities, which is 3/10 + 1/10 = 4/10 = 2/5.

Answer: A

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion